OpenStudy (anonymous):

The value of the discriminant for the equation below is 33. 2x^2+5x-1=0 What does the value of the discriminant mean for this equation?

OpenStudy (anonymous):

Right. But since the discriminant is 33 does this mean that the equation has 1 real, rational root, 2 real rational roots, 1 reall irrational root or 2 real irrational roots?

OpenStudy (anonymous):

not to be picky but the discriminant is \(b^2-4ac\) and not \(\sqrt{b^2-4ac}\)

OpenStudy (anonymous):

\[2x^2+5x-1=0\] discriminant is 33 and since \(33>0\) you have two real roots. also since \(33\) is not a perfect square, that means \(\sqrt{33}\) is irrational

OpenStudy (anonymous):

there is only one real irrational root, then, right? i'm confused.

OpenStudy (anonymous):

thank you